A 99% confidence interval for  is given by:  

The multiplier is __________ .

A. 2.576

B. 1.96

C. None of the other choices represent a suitable response.

D. 1.282

E. 3
Answer #1

Answer: A. 2.576

Explanation:

We are given

Confidence level = c = 99% = 0.99

α = 1 - 0.99 = 0.01

α/2 = 0.01/2 = 0.005

1 - α/2 = 1 - 0.005 = 0.995

Z(1 - α/2) = Z0.995 = 2.576

(by using z-table)

Answer:

A. 2.576